### v1

# 2TRIUMF, 4004 Wesbrook Mall, Vancouver, BC V6T 2A3 Canada

Abstract
By extending our analysis and simulations of our 37K $\beta$-decay data set to allow the $\beta$ asymmetry with respect to nuclear spin to vary with $\beta$ energy $E_{\beta}$, we have gained sensitivity to new physics that depends on a helicity factor for the $\beta$, $m_{\beta}/E_{\beta}$. In particular, we constrain Lorentz scalar and tensor quark-lepton interaction strengths at a sensitivity complementary to the similar Fierz interference term in neutron $\beta$ decay. Our result for that new physics is $b_{F}$ = -0.0002 $\pm$ 0.024 (stat) $\pm$ 0.039 (syst), consistent with the standard model electroweak interaction value $b_{F}=0$. We consider presently achieved complementarity to $\beta$-decay and particle physics experiments, along with projectable technical improvements to our method.
preprint: APS/123-QED
I Introduction
The standard model electroweak interaction always makes left-handed helicity leptons paired with right-handed helicity antileptons. In a previous publication [4], we constrained new physics coupling to flipped helicity for both the $\beta$ and the $\nu$, i.e. quark-lepton interactions Lorentz transforming like vector (V) + axial vector (A) currents, rather than the weak interaction’s V-A Lorentz current structure. Our observable, the asymmetry of $\beta$ emission with respect to nuclear spin Aβ, was predicted to be independent of total $\beta$ energy Eβ either for V-A or V+A.
Here we extend simulations and analysis techniques to allow for different new physics that couples a $\nu$ with weak interaction helicity to a $\beta^{+}$ with the same sign of helicity as the $\nu$, which mutiplies the result by a helicity factor $m_{\beta}/E_{\beta}$ needed to produce that $\beta$ with opposite helicity. Having noted in Ref. [4] that one dominant systematic, $\beta$ scattering off surfaces, strongly depends on $E_{\beta}$ as it approaches $m_{\beta}$, we find that the energy dependence complicates the simulations and analysis considerably.
II Theory and Methods
New physics that couples outgoing leptons and antileptons with same helicity– rather than opposite helicity as in the standard model weak interaction– produces a decay contribution dependent on a helicity factor of the massive $\beta$, $m_{\beta}/E_{\beta}$, the Fierz interference term between the new physics and the standard model matrix elements.
The decay rate expression then adds the Fierz term $b$:
$\frac{dW}{dE}=F(E,Z)pEp_{\nu}E_{\nu}(1+bm/E+\hat{I}\cdot(A_{\beta}\frac{\vec{p_{\beta}}}{E})),$ (1)
The normalized coefficient of $\cos{\theta}$, where $\theta$ is the angle between the $\beta$ and the initial nuclear polarization, becomes a function of the $\beta$ energy:
$A_{\beta}[E]=A_{\beta}/(1+mb/E)\approx A_{\beta}(1-mb/E)$ (2)
We write the approximate expression to emphasize that sensitivity to $b$ in any given decay scales with $A_{\beta}$. The value of $A_{\beta}$ in 37K is reasonably large compared to other well-measured cases like the neutron and 19Ne, making 37K decay more sensitive to $b$ by that simple figure of meri.
We now specialize slightly to the Lee-Yang Lagrangian for electron-nucleon interactions, and call its Fierz interference term explicitly $b_{F}$. It has a Lorentz scalar structure interfering with the standard model vector, and a tensor interfering with axial vector [1]:
$b_{F}=\frac{\pm\gamma[(C_{S}+C_{S}^{\prime})/C_{V}+\rho^{2}(C_{T}+C_{T}^{\prime})/C_{A}]}{1+\rho^{2}+(C_{S}^{2}+C_{S}^{\prime 2}+C_{T}^{2}+C_{T}^{\prime 2})/2}$ (3)
where $\gamma$$\sqrt{1-Z^{2}\alpha^{2}}$. For isobaric mirror decay
$A_{\beta}=\frac{\frac{\pm\rho^{2}}{J+1}(C_{T}C_{T}^{\prime}-C_{A}C_{A}^{\prime})}{1+\rho^{2}+((C_{S}^{2}+C_{S}^{\prime 2})+(C_{T}^{2}+C_{T}^{\prime 2})/2)}$ (4)
$+\frac{\rho\sqrt{\frac{J}{J+1}}(C_{S}C_{T}^{\prime}+C_{S}^{\prime}C_{T}-C_{V}C_{A}^{\prime}-C_{V}^{\prime}C_{A})}{1+\rho^{2}+(C_{S}^{2}+C_{S}^{\prime 2}+C_{T}^{2}+C_{T}^{\prime 2})/2}$ (5)
where $\rho$ is the standard generalization for nuclei $\rho=\frac{g_{A}}{g_{V}}\frac{M_{GT}}{M_{F}}$ determined from an average Fermi strength value along with the measured 37K $fT$ value [19]. The $\pm$ signs are for $\beta^{\mp}$ decay.
We note that the Lee-Yang Lagrangian assumes zero-range interactions, i.e. interactions carried by bosons with mass much greater than the $\beta$ decay energy release, so that terms from expanding the boson propagator are negligible. It otherwise characterizes completely general types of Lorentz structures of quark-lepton interactions.
[…]
A 2nd-class term in the axial vector lepton-nucleon current, termed the induced tensor term $\frac{g_{T}}{m}\sigma_{\mu\nu}k_{\nu}$, has different dependence on $E_{\beta}$ than the electron-quark Lorentz tensor interaction [6, 21]. It can be generated by non-standard-model meson exchange and nucleon-nucleon contact interactions [21, 22, 23], and there are tight constraints from tailored nuclear $\beta$ decay observables, in particul the $\beta$-alignment correlation which vanishes in Gamow-Teller decay [12]. In this parameterization of 2nd-class tensor interactions, one term does contribute equally to neutron and nuclear $\beta$ decay, while two other complicated nuclear matrix elements can vary with any nucleus, only one of which vanishes in isobaric mirror decays [22, 23]. Relating constraints on $C_{T}$ is nontrivial, so we don’t comment further below.
A number of experimental methods worldwide measure small changes in the $\beta$ energy spectrum itself. Instead, we focus on the asymmetry of emission of $\beta$’s with respect to the initial spin. Since the asymmetry is not changing much with energy– mainly with v/cv/c– this minimizes dependence on systematics like the $\beta$ detector lineshape response.
II.1 Recoil-order corrections
We treat in our standard model description Coulomb and 1st and 2nd-order recoil corrections according to Holstein [6]. Several recoil-order terms in isobaric mirror decay are given by the conserved vector current hypothesis– e.g. weak magnetism and 2nd-order in recoil electric quadrupole moment analog are given by their respective nuclear moment differences in parent and progeny, while the first-class induced tensor term vanishes. The $\mu$ moment of 37K, and thus the weak magnetism term, are small, while the $\beta$ energy endpoint near 5 MeV makes the 2nd-order corrections similar in magnitude. We found that corrections can be combined to add $\approx$ -0.0028 $E_{\beta}$$E_{0}$ to $A_{\beta}$ [4]. A number of careful corrections to Ref. [6] were shown small in our case [10].
III Experiment
III.1 TRIUMF Neutral Atom Trap
Fig. 1 is a side view of the detection apparatus of TRIUMF’s Neutral Atom Trap for $\beta$ decay (TRINAT). Not shown is the collection trap from a vapor cell cube, nor the push beams between traps [2].
Figure 1: TRINAT during the optical pumping time. Shown are $\beta$ telescopes (double-sided silicon strip detector DSSD $\Delta E$ backed by plastic scintillator), mirrors for optical pumping light and its beams, magnetic field coils, electric field electrodes, and microchannel plates for electron and ion detection. Distance between trap cloud and ion MCP is 9.7 cm.
Using 8x107/s mass-separated 37K delivered from the TRIUMF/ISAC isotope separator on-line facility, we trapped on average 10,000 37K atoms during the data-taking time. ISAC, by using a microstructured TiC target, has recently demonstrated 4 times more yield.
III.2 Polarization by optical pumping
We refer the reader to our in situ measurement of the nuclear polarization by atomic methods, Ref. [3] and the Supplementary Material of Ref. [4].
During the polarization time, we switch the magnetic field from the trap’s quadrupole to a uniform 2 Gauss field pointed up, and apply circularly polarized light along the quantization axis. Once we start the optical cycle, atoms increase spin to maximum, then stop absorbing the S1/2 to P1/2 light.
By atomic physics probes of this process independent of the $\beta$ decays, we measured nuclear polarization $I_{z}/I=0.991\pm 0.001$, which contributes negligibly to our uncertainty budget.
III.3 Geometry and Detectors
An electric field collects 37Ar ions produced in 37K $\beta^{+}$ decay, along with photoions from the 355 nm laser in Fig. 1, to an MCP with 78 mm active diameter located 9.7 cm away. By comparing trajectories with a detailed finite element calculation, we found our simulations could assume a uniform 150 V/cm field. The photoion information hleps determine the cloud polarization, center in space, and spatial distribution. The 37Ar ion info is not used in this paper’s analysis.
Decay by $\beta^{+}$ naturally feeds the negative Ar-1 ion. We constrained contributions from the known Ar-1 $\tau$=250 ns metastable by the lack of a significant tail on the timing spectrum between shakeoff electron and $\beta$: nevertheless, to eliminate possible bias on $A_{\beta}$ from loss of detection efficiency from the delayed shakeoff emission and its dependence on the recoil energy and direction, we don’t impose a vertical position cut on the shakeoff electron detector [4].
We used 0.30 mm thick double-sided silicon strip detectors as Ref. [4], requiring both X and Y strips above energy threshold and similar calibrated energy deposited. Our $E$ detectors are plastic 3.5x9 cm scintillators shown, characterized further in Ref. [5].
III.4 Data and interpretation
We found it important in Ref. [4] to remove backgrounds in our geometry from decays of untrapped atoms by requiring the $\beta^{+}$ to be in coincidence with shakeoff atomic electrons (SOE’s). To improve our knowledge of the asymmetry at low $E_{\beta}$ to study $b_{F}$, we found the relative timing to be even more important. So we removed a residual timing walk in the scintillator, produced at low $E_{\beta}$ pulse height, which had not been removed by the constant fraction discriminator. The resulting improved timing (Fig. 2) nearly removes a small background from untrapped atoms on surfaces that is more apparent at low $E_{\beta}$. The simulated background is scaled to a tail of the timing spectrum, as we do not know what fraction of untrapped atoms sticks long enough to decay. This background was one of our larger systematics in Ref. [4], while with the improved timing the correction from the simulation shown is nearly negligible and introduces small uncertainty. Also, we can now understand and model well the measured timing signal from the known cloud size using a combination of 9% very low-energy electrons– probably Ar- autoionization– 77% 4S electron shakeoff characterized by their 4.5 eV binding energy, and 14% 3P electron shakeoff characterized by their 18 eV binding energy. The SOE energy spectra use the sudden approximation calculations with hydrogenic wavefunctions by Levinger [11].
Figure 2: Improved shakeoff electron (SOE) timing wrt the $\beta$ compared to Fig. 4 of Ref. [4], and the resulting smaller background. The data timing cut is shown. The summed simulation reproduces the data well in the region of interest. 10% of real coincidences are from electrons not firing eMCP, scattering, and returning in the E field 14 ns later– note this late 10% of the prompt $\beta$-$\gamma$ coincidences do not fall in the ROI.
III.4.1 Simulations
We show our GEANT4 simulation against half the data in Fig. 3. A superratio observable combining up and down $\beta$ detectors and up and down nuclear spin polarization measures the $\beta$ asymmetry. The inflection region between 1.5 and 2 MeV is critical, and the improved fit here removes a small systematic apparent in our previous result Ref. [4]. The fit threshold is the same as in Ref. [4], determined by the Compton edge from 511 keV annihilation radiation in our plastic scintillator $E$ detector, together with energy loss in our optical pumping mirrors and our DSSD $\Delta E$ detector. The other half of the data has a slightly different energy calibration, so we do not choose to show a combined plot.
[…]
III.5 Extraction of $A_{\beta}[E_{\beta}]$
We extract $A_{\beta}$ and $b_{F}$ from this info. We calculate $\chi^{2}$ from the difference between the superratio of all data and simulations with varied $A_{\beta}$ and $b_{F}$.
[…]
To extract our two degrees of freedom, we then look for this total $\chi^{2}$ to increase from its minimum by 2.3 when varying $A_{\beta}$ and $b_{F}$ [13]. The result, shown in Fig. 4, is $b_{F}=-0.0002\pm 0.024$ (statistical). We consider systematics in the next section.
Figure 4: A $\chi^{2}$ map of all experimental data compared to a simulated parameter space of $A_{\beta}$ and $b$, showing in solid lines the 1 $\sigma$ 68% confidence limits from a change in $\chi^{2}$ of 2.3 from the minimum. Dashed lines are 90% CL. The tighter, black ellipses add the $chi^{2}$ contribution from the predicted center value of $A_{\beta}$, which determines $b_{F}$ of Eq. 3 from electron-quark Lorentz scalar and tensor intereactions. The larger, red ellipses determine the more general measurement of $b$ from Eq. 1 without the FtFt constraint.
III.5.1 Peripheral quantities
For completeness, we consider two other extremes to extract information from our data set.
As a check, by forcing $b_{F}$=0, i.e. introducing only the same energy-independent physics allowed by our Ref. [4], we deduce $A_{\beta}$= -0.5712 $\pm$ 0.013 (stat). Since forcing $b_{F}$=0 produces a 1-parameter fit, we read off the change in $\chi^{2}$ by 1, not the 2.3 change for the 2-parameter fits. This would imply that our reanalysis in this paper, primarily the time walk experimental correction of Fig. 2, would alter our previous answer for the $\beta$ asymmetry with no extra energy dependence from $A_{\beta}=-0.5707(13)(13)$ to –0.5713(13)(10). The smaller systematic uncertainty is from the near-elimination of the background from untrapped atoms. The centroid change is about one-third of our total uncertainty, so the upper limits on right-handed currents in Ref. [4] and the extraction of $V_{ud}$ are essentially unchanged, and we do not comment further.
Although the Lee-Yang Lagrangian is the only model we are aware of for $b$, we also consider a more general situation where $b$ and $A_{\beta}$ are allowed to vary independently without the prediction of $A_{\beta}$ from the 37K $fT$ value. E.g, this would happen if V+A, S, and T currents were all allowed to vary independently. It should be intuitive that there are large correlations between $A_{\beta}$ and $b$ in Fig. 3, so the result inflates the uncertainty on both: $b$ = 0.033 $\pm$ 0.084 (stat) $\pm$ 0.0039 (syst) and $A_{\beta}=-0.5738\pm 0.0082{\rm(stat)}\pm 0.0041{\rm(syst)}$. We re-state for emphasis that these are not primary results of this paper.
III.6 Systematic uncertainties
We show systematic uncertainties in Table I. Most are determined by testing whether a nonzero $b$ is introduced when a parameter is varied in the GEANT4 simulation. We describe a few critical differences from our earlier analysis in Ref. [4].
Beta Scattering The largest systematic uncertainty is from $\beta$ scattering, addressed carefully in our Ref. [4]. Our Supplementary Material showed our own DSSD double-firing events, which provided a solid benchmark to our version of GEANT4 simulation including different low-energy physics lists. This scattering increases at our lower $E_{\beta}$, potentially introducing false asymmetries, and we find this determines our largest systematic. We project that by using low-Z materials we can decrease this by an order of magnitude.
[…]
Low-$E_{\beta}$ lineshape and energy threshold The measured asymmetry reflects the standard model’s dependence on helicity, the non-unity velocity of the $\beta$ at our lower energies apparent in Fig. 3. So experimental features changing the energy dependence, such as the natural tail in scintillator response from bremsstrahlung and backscatter, introduce additional uncertainty in the energy dependence of the asymmetry. Lowering the $E_{\beta}$ threshold with deomonstrated much thinner mirrors and in-progress $\Delta E$ would also improve our statistical sensitivity to the $m_{\beta}/E_{\beta}$ term by a factor of two.
Source Present Future
$b$ $b$
$\beta$ Scattering 0.031 0.003
Mirror Thickness 0.013 0.001
DSSD Thickness 0.013 0.001
DSSD Detection Radius 0.006 0.001
DSSD signal/noise 0.006 0.001
Low-$E_{\beta}$ lineshape 0.008 0.008
DSSD XY Energy Agreement 0.005 0.001
DSSD $E$ threshold 0.005 0.001
Scintillator Threshold 0.004 0.001
Scintillator Calibration 0.003 0.003
Atomic Cloud 0.002 0.002
Background 0.004 0.004
Be Foil Thickness 0.004 0.004
Total Systematics 0.039 0.011
Table 1: Systematic uncertainty budget for the two-parameter analysis for $b$ (which also applies to $b_{F}$) and $A_{\beta}$. All uncertainties are believed to be uncorrelated, and are added in quadrature. See text for future projections
IV Results
Our result in the Lee-Yang scalar-tensor Lagrangian of quark-lepton non-SM Lorentz current is then
$b_{F}$ = -0.0002 $\pm$ 0.024 (stat) $\pm$ 0.039 (syst),
We explore complementarity of this our result to other experiments next.
IV.1 Sensitivity to new physics
We show a 90% CL (i.e. 1.65 $\sigma$) exclusion plot of Lorentz scalar vs. tensor lepton-quark couplings in Fig. 5, assuming no coupling to non-standard model right-handed neutrinos. There exist complete reviews setting constraints on the Lee-Yang Lagrangian parameters [14], and separately reviews on 2nd-class currents [21, 12]. To allow for the possibility of non-universal interactions, here we show ten $\beta$ decay and other particle physics constraints, without review or averaging. Our goal is to show complementarity in situations with less universal physics models than the Lee-Yang Lagrangian.
Figure 5: Exclusion plot of Lorentz scalar and tensor lepton-quark couplings, 90% CL. Consistency between neutron $\beta$ decay, nuclear $\beta$ decay, and particle physics requires a non-universal interaction with scale less than TeV. See details and references in text.
Our 37K results could be said to complement the neutron results on $C_{S}$ because our smaller Gamow-Teller strength weighs $C_{S}$ more than $C_{T}$, yet our own group’s direct constraints on $C_{S}$ from 38mK decay [26] are considerably tighter, shown here averaged with the 32Ar experiment with similar sensitivity [24, 25].
We show three constraints from particle physics with computable sensitivity to the Lee-Yang Lagrangian and no sensitivity to 2nd-class currents. Strong limits are set on $C_{T}=C_{T}^{\prime}$ from $\pi\rightarrow e\nu\gamma$ (blue cyan band) [39, 49], considerably stronger than the combined limit from 6He  [40, 41, 42, 43], 8B, and 8Li aβν [44, 45], as well as the combined limit from three Gamow-Teller $A_{\beta}$ measurements with similar sensitivity due to relatively large $\langle m_{\beta}/E_{\beta}\rangle$ [14, 46, 47, 48].
An effective field theory relating $\beta$ decay to an isospin rotation of LHC p+p $\rightarrow$ electron + missing transverse mass $m_{\perp}$ (i.e. a $\nu$) tightly constrains new physics with $m_{\perp}\geq$ 1.5 TeV from $\sqrt{s}$=8 TeV data (open blue circle) [27], and even more tightly from $\sqrt{s}$=13 TeV data if $m_{\perp}\geq 2.2~{\rm TeV}$ [14, 15, 16, 29, 30].
By adding a W in loop diagrams, limits on pseudoscalar interactions from $\pi\rightarrow e\nu$ imply a tight constraint $0.0012\leq C_{S}\leq 0.00027$ at 90% CL (blue solid vertical lines) for a 200 GeV new boson, with limits becoming weaker by an order of magnitude for various models (a new boson at the Z mass, scalars coupling purely to right-handed $\nu$’s, or purely imaginary couplings) [18]. Both the LHC channel and the $\pi\rightarrow e\nu$ constraints have multiple EFT operators contributing [14, 18] which could in principle cancel in an explicit model of microscopic physics.
The other tight indirect constraints on $C_{S}$, shown in the green vertical band, are from the consistency of $0^{+}\rightarrow 0^{+}$ $Ft[E_{\beta}]$ values in a 2020 global survey [31]. Yet allowing $b_{F}$ to float instead of fixing it to zero would inflate the uncertainty in $V_{ud}$ by factors of two, motivating independent determinations of $b_{F}$ with similar sensitivity. Recent theory developments may relax or improve the sensitivity: re-evaluation of the lepton phase space $f$ summarized in Ref. [50] with better calculation of a radiative correction in 10C decay [51]; continuing theoretical challenges to support existing isospin-breaking calculations, including a possible need for first-principles inclusion of strong interaction isospin breaking [52] rather than fitting the Nolen-Schiffer anomaly [31]. Note that a fit to Ref. [31] of the 2nd-class induced scalar contribution to an electron-nucleon vector current parameterized as $\frac{e}{A}\frac{m_{\beta}}{m_{\rm nucleon}}\frac{m_{\beta}}{E_{\beta}}$ [54, 20], produces $e=-4\pm 32$.
Two experiments are shown for neutron $\beta$ decay. Measurement of $\beta$ decay asymmetry and spectrum shape has competitive sensitivity to $b_{F}$ consistent with zero (solid black lines) [32]. An analysis including the recoil energy spectrum of protons and Ref. [32] data together, but floating $g_{A}/g_{V}$ to a different result from either, asserts a 2.8 $\sigma$ nonzero $b_{F}$ result (the dashed black line) [33]. By inspection, the positive result of Ref. [33] can be reconciled with the tight constraint on Lorentz tensor lepton-quark interactions from $\pi\rightarrow e\nu\gamma$ decay [39] with a nonzero and negative $C_{S}$ of -0.05. There are by inspection varying degrees of tension with the above-mentioned Lorentz electron-quark scalar current constraints [26, 24, 25, 31], with isobaric mirror decays like the present 37K work, with 19Ne (from Ref. [35] which used its $fT$ measurement together with $A_{\beta}[E_{\beta}]$ [34], both under revision [36]), and with the relative helicity of $\beta$’s in Fermi vs. Gamow-Teller decay (the cyan band denoted $P_{F}/P_{GT}$[38]. Neverthless, we note that interpreting the $C_{S}\approx$ -0.05 region as an induced scalar for neutron decay $e\approx$ 31 alleviates tension with all other nuclear $\beta$ decay because of its $1/A$ dependence.
We see from Fig. 5 that to be complementary to other physics, we would have to improve our 37K $b_{F}$ result to compete with the best isospin mirror nucleus results in 19Ne [35, 34, 36]; measurements of $a_{\beta\nu}$ in 21Na [37] would need similar improvement. We hope demonstrated and in-progress improvements in the Experiment suggestions will let us achieve the factor of five needed.
V Conclusions
We have used our $\beta$ asymmetry data from the decay of 37K to constrain new physics that depends on the $\beta$ energy, measuring the Fierz interference term from quark-lepton Lorentz scalar and vector couplings $b_{F}$ = -0.0002 $\pm$ 0.024 (stat) $\pm$ 0.039 (syst), We provide constraints on a combination of Lorentz scalar and tensor lepton-quark couplings. Our present results using our laser-cooled atoms are dominated by a combination of statistics and solvable $\beta$ scattering, and we hope to improve sensitivity by a factor of about five to enable complementarity with experiments indicating possible existence of non-standard model physics.
Acknowledgements.
We acknowledge TRIUMF/ISAC staff, in particular for TiC target preparation. Supported by the Natural Sciences and Engineering Research Council of Canada, and the U.S. Department of Energy, Office of Science, Office of Nuclear Physics under Awards No. DE-FG03- 93ER40773 and No. DE-FG02-11ER41747. TRIUMF receives federal funding via a contribution agreement through the National Research Council of Canada.

### v4

@@
-2TRIUMF, 4004 Wesbrook Mall, Vancouver, BC V6T 2A3 Canada
+
 Abstract
@@
 In particular, we constrain Lorentz scalar and tensor quark-lepton interaction strengths at a sensitivity complementary to the similar Fierz interference term in neutron $\beta$ decay.
-Our result for that new physics is $b_{F}$ = -0.0002 $\pm$ 0.024 (stat) $\pm$ 0.039 (syst), consistent with the standard model electroweak interaction value $b_{F}=0$.
+Our result for that new physics is a Fierz interference term $b$ = 0.033 $\pm$ 0.084 (stat) $\pm$ 0.039 (syst), consistent with the standard model electroweak interaction value $b=0$.
 We consider presently achieved complementarity to $\beta$-decay and particle physics experiments, along with projectable technical improvements to our method.
@@
 Our observable, the asymmetry of $\beta$ emission with respect to nuclear spin Aβ, was predicted to be independent of total $\beta$ energy Eβ either for V-A or V+A.
-Here we extend simulations and analysis techniques to allow for different new physics that couples a $\nu$ with weak interaction helicity to a $\beta^{+}$ with the same sign of helicity as the $\nu$, which mutiplies the result by a helicity factor $m_{\beta}/E_{\beta}$ needed to produce that $\beta$ with opposite helicity.
+Here we extend simulations and analysis techniques to allow for different new physics that couples a $\nu$ with weak interaction helicity to a $\beta^{+}$ with the same sign of helicity as the $\nu$, which multiplies the result by a helicity factor $m_{\beta}/E_{\beta}$ needed to produce that $\beta$ with opposite helicity.
 Having noted in Ref.
@@
 The normalized coefficient of $\cos{\theta}$, where $\theta$ is the angle between the $\beta$ and the initial nuclear polarization, becomes a function of the $\beta$ energy:
-$A_{\beta}[E]=A_{\beta}/(1+mb/E)\approx A_{\beta}(1-mb/E)$ (2)
+$A_{\beta}[E]=A_{\beta}/(1+bm/E)\approx A_{\beta}(1-bm/E)$ (2)
 We write the approximate expression to emphasize that sensitivity to $b$ in any given decay scales with $A_{\beta}$.
-The value of $A_{\beta}$ in 37K is reasonably large compared to other well-measured cases like the neutron and 19Ne, making 37K decay more sensitive to $b$ by that simple figure of meri.
-We now specialize slightly to the Lee-Yang Lagrangian for electron-nucleon interactions, and call its Fierz interference term explicitly $b_{F}$.
-It has a Lorentz scalar structure interfering with the standard model vector, and a tensor interfering with axial vector [1]:
-$b_{F}=\frac{\pm\gamma[(C_{S}+C_{S}^{\prime})/C_{V}+\rho^{2}(C_{T}+C_{T}^{\prime})/C_{A}]}{1+\rho^{2}+(C_{S}^{2}+C_{S}^{\prime 2}+C_{T}^{2}+C_{T}^{\prime 2})/2}$ (3)
+The value of $A_{\beta}$ in 37K is reasonably large compared to other well-measured cases like the neutron and 19Ne, making 37K decay more sensitive to $b$ by that simple figure of merit.
+In the Lee-Yang Lagrangian, $b$ is produced by a Lorentz scalar interaction interfering with the standard model Lorentz vector interaction, and a Lorentz tensor interaction interfering with a Lorentz axial vector interaction [1]:
+$b=\frac{\pm\gamma[(C_{S}+C_{S}^{\prime})+\rho^{2}(C_{T}+C_{T}^{\prime})]}{1+\rho^{2}+(C_{S}^{2}+C_{S}^{\prime 2}+\rho^{2}(C_{T}^{2}+C_{T}^{\prime 2}))/2}$ (3)
 where $\gamma$$\sqrt{1-Z^{2}\alpha^{2}}$. For isobaric mirror decay
-$A_{\beta}=\frac{\frac{\pm\rho^{2}}{J+1}(C_{T}C_{T}^{\prime}-C_{A}C_{A}^{\prime})}{1+\rho^{2}+((C_{S}^{2}+C_{S}^{\prime 2})+(C_{T}^{2}+C_{T}^{\prime 2})/2)}$ (4)
-$+\frac{\rho\sqrt{\frac{J}{J+1}}(C_{S}C_{T}^{\prime}+C_{S}^{\prime}C_{T}-C_{V}C_{A}^{\prime}-C_{V}^{\prime}C_{A})}{1+\rho^{2}+(C_{S}^{2}+C_{S}^{\prime 2}+C_{T}^{2}+C_{T}^{\prime 2})/2}$ (5)
-where $\rho$ is the standard generalization for nuclei $\rho=\frac{g_{A}}{g_{V}}\frac{M_{GT}}{M_{F}}$ determined from an average Fermi strength value along with the measured 37K $fT$ value [19].
+$A_{\beta}=\frac{\frac{\pm\rho^{2}}{J+1}(C_{T}C_{T}^{\prime}-1)}{1+\rho^{2}+(C_{S}^{2}+C_{S}^{\prime 2}+\rho^{2}(C_{T}^{2}+C_{T}^{\prime 2}))/2}$
+$+\frac{\rho\sqrt{\frac{J}{J+1}}(C_{S}C_{T}^{\prime}+C_{S}^{\prime}C_{T}-2)}{1+\rho^{2}+(C_{S}^{2}+C_{S}^{\prime 2}+\rho^{2}(C_{T}^{2}+C_{T}^{\prime 2}))/2}$ (4)
+where $\rho=\frac{g_{A}}{g_{V}}\frac{M_{GT}}{M_{F}}$ is determined by us below from our collaboration’s measured 37K $fT$ value [19], together with an average $M_{F}$ from $0^{+}\rightarrow 0^{+}$ decays [32].
 The $\pm$ signs are for $\beta^{\mp}$ decay.
@@
 […]
-A 2nd-class term in the axial vector lepton-nucleon current, termed the induced tensor term $\frac{g_{T}}{m}\sigma_{\mu\nu}k_{\nu}$, has different dependence on $E_{\beta}$ than the electron-quark Lorentz tensor interaction [6, 21].
-It can be generated by non-standard-model meson exchange and nucleon-nucleon contact interactions [21, 22, 23], and there are tight constraints from tailored nuclear $\beta$ decay observables, in particul the $\beta$-alignment correlation which vanishes in Gamow-Teller decay [12].
+A 2nd-class term in the axial vector lepton-nucleon current, termed the induced tensor term $\frac{g_{T}}{m_{\rm nucleon}}\sigma_{\mu\nu}k_{\nu}$, has different dependence on $E_{\beta}$ than the electron-quark Lorentz tensor interaction [6, 21].
+It can be generated by non-standard-model meson exchange and nucleon-nucleon contact interactions [21, 22, 23], and there are tight constraints from tailored nuclear $\beta$ decay observables, in particular the $\beta$-alignment correlation which vanishes in Gamow-Teller decay [12].
 In this parameterization of 2nd-class tensor interactions, one term does contribute equally to neutron and nuclear $\beta$ decay, while two other complicated nuclear matrix elements can vary with any nucleus, only one of which vanishes in isobaric mirror decays [22, 23].
-Relating constraints on $C_{T}$ is nontrivial, so we don’t comment further below.
+Relating constraints to those on $C_{T}$ is nontrivial, so we don’t comment further below.
 A number of experimental methods worldwide measure small changes in the $\beta$ energy spectrum itself.
@@
 By comparing trajectories with a detailed finite element calculation, we found our simulations could assume a uniform 150 V/cm field.
-The photoion information hleps determine the cloud polarization, center in space, and spatial distribution.
+The photoion information helps determine the cloud polarization, center in space, and spatial distribution.
 The 37Ar ion info is not used in this paper’s analysis.
 Decay by $\beta^{+}$ naturally feeds the negative Ar-1 ion.
-We constrained contributions from the known Ar-1 $\tau$=250 ns metastable by the lack of a significant tail on the timing spectrum between shakeoff electron and $\beta$: nevertheless, to eliminate possible bias on $A_{\beta}$ from loss of detection efficiency from the delayed shakeoff emission and its dependence on the recoil energy and direction, we don’t impose a vertical position cut on the shakeoff electron detector [4].
+We constrained contributions from the known Ar-1 $\tau$=250 ns metastable state by the lack of a significant tail on the timing spectrum between shakeoff electron and $\beta$: nevertheless, to eliminate possible bias on $A_{\beta}$ from loss of detection efficiency from the delayed shakeoff emission and its dependence on the recoil energy and direction, we don’t impose a vertical position cut on the shakeoff electron detector [4].
 We used 0.30 mm thick double-sided silicon strip detectors as Ref.
@@
 [4] to remove backgrounds in our geometry from decays of untrapped atoms by requiring the $\beta^{+}$ to be in coincidence with shakeoff atomic electrons (SOE’s).
-To improve our knowledge of the asymmetry at low $E_{\beta}$ to study $b_{F}$, we found the relative timing to be even more important.
+To improve our knowledge of the asymmetry at low $E_{\beta}$ to study $b$, we found the relative timing to be even more important.
 So we removed a residual timing walk in the scintillator, produced at low $E_{\beta}$ pulse height, which had not been removed by the constant fraction discriminator.
@@
 III.5 Extraction of $A_{\beta}[E_{\beta}]$
-We extract $A_{\beta}$ and $b_{F}$ from this info. We calculate $\chi^{2}$ from the difference between the superratio of all data and simulations with varied $A_{\beta}$ and $b_{F}$.
-[…]
-To extract our two degrees of freedom, we then look for this total $\chi^{2}$ to increase from its minimum by 2.3 when varying $A_{\beta}$ and $b_{F}$ [13].
-The result, shown in Fig.
-4, is $b_{F}=-0.0002\pm 0.024$ (statistical).
-We consider systematics in the next section.
+We extract $A_{\beta}$ and $b$ from this info.
+We calculate $\chi^{2}$ from the difference between the superratio of all data and simulations with varied $A_{\beta}$ and $b$.
+To extract our two degrees of freedom, we then look for this total $\chi^{2}$ to increase from its minimum by 2.3 when varying $A_{\beta}$ and $b$ [13].
+Allowing $b$ and $A_{\beta}$ to vary independently produces our main result, $b$ = 0.033 $\pm$ 0.084 (stat) $\pm$ 0.0039 (syst) and $A_{\beta}=-0.5738\pm 0.0082{\rm(stat)}\pm 0.0041{\rm(syst)}$.
+The linear correlation between $A_{\beta}$ and $b$ reflected in Fig.
+4 can be expressed by $A_{\beta}=-0.5707-0.090*b$.
+We find more sensitivity to new physics by adding information.
+We also show the standard model prediction for $A_{\beta}=-0.5706$ in Fig.
+4, given $\rho$ and its uncertainty from the measured 37K $Ft$ of our collaboration [19], and taking $M_{F}$ from the measured $0^{+}\rightarrow 0^{+}$ average $Ft_{0}$ value [32].
+In Fig.
+4 we are not considering the variation in $Ft_{0}$ with $C_{S}+C^{\prime}_{S}$, as we wish to show 37K complementarity to such observables below.
+We further show $\chi^{2}$ contours as black ellipses in Fig.
+4 including the contribution from the experimental uncertainty on 37K $Ft$.
+When Lorentz scalar and tensor interactions are considered below, we correlate $b$ variations with changes in $A_{\beta}$ by Eqs.
+3-4, and plot the resulting combined constraint on those couplings in Fig.
+5.
 Figure 4: A $\chi^{2}$ map of all experimental data compared to a simulated parameter space of $A_{\beta}$ and $b$, showing in solid lines the 1 $\sigma$ 68% confidence limits from a change in $\chi^{2}$ of 2.3 from the minimum.
 Dashed lines are 90% CL.
-The tighter, black ellipses add the $chi^{2}$ contribution from the predicted center value of $A_{\beta}$, which determines $b_{F}$ of Eq.
-3 from electron-quark Lorentz scalar and tensor intereactions.
 The larger, red ellipses determine the more general measurement of $b$ from Eq.
-1 without the FtFt constraint.
-III.5.1 Peripheral quantities
-For completeness, we consider two other extremes to extract information from our data set.
-As a check, by forcing $b_{F}$=0, i.e.
+1.
+The tighter, black ellipses add the $chi^{2}$ contribution from the predicted center value of $A_{\beta}$ from FtFt information (see text), which determines $C_{S}+C^{\prime}_{S}$ vs.
+$C_{T}+C^{\prime}_{T}$ in Fig.
+5 below.
+III.5.1 Check of previous result
+As a check, by forcing bb=0, i.e.
 introducing only the same energy-independent physics allowed by our Ref.
 [4], we deduce $A_{\beta}$= -0.5712 $\pm$ 0.013 (stat).
-Since forcing $b_{F}$=0 produces a 1-parameter fit, we read off the change in $\chi^{2}$ by 1, not the 2.3 change for the 2-parameter fits.
-This would imply that our reanalysis in this paper, primarily the time walk experimental correction of Fig.
-2, would alter our previous answer for the $\beta$ asymmetry with no extra energy dependence from $A_{\beta}=-0.5707(13)(13)$ to –0.5713(13)(10).
+Since forcing bb=0 produces a 1-parameter fit, we read off the change in $\chi^{2}$ by 1, not the 2.3 change for the 2-parameter fits.
+This deductions implies that our reanalysis in this paper, primarily the time walk experimental correction of Fig.
+2, alters our previous answer for the $\beta$ asymmetry with no extra energy dependence from $A_{\beta}=-0.5707(13)(13)$ to -0.5712(13)(10).
 The smaller systematic uncertainty is from the near-elimination of the background from untrapped atoms.
@@
 [4] and the extraction of $V_{ud}$ are essentially unchanged, and we do not comment further.
-Although the Lee-Yang Lagrangian is the only model we are aware of for $b$, we also consider a more general situation where $b$ and $A_{\beta}$ are allowed to vary independently without the prediction of $A_{\beta}$ from the 37K $fT$ value.
-E.g, this would happen if V+A, S, and T currents were all allowed to vary independently.
-It should be intuitive that there are large correlations between $A_{\beta}$ and $b$ in Fig.
-3, so the result inflates the uncertainty on both: $b$ = 0.033 $\pm$ 0.084 (stat) $\pm$ 0.0039 (syst) and $A_{\beta}=-0.5738\pm 0.0082{\rm(stat)}\pm 0.0041{\rm(syst)}$.
-We re-state for emphasis that these are not primary results of this paper.
 III.6 Systematic uncertainties
@@
 Total Systematics 0.039 0.011
-Table 1: Systematic uncertainty budget for the two-parameter analysis for $b$ (which also applies to $b_{F}$) and $A_{\beta}$.
-All uncertainties are believed to be uncorrelated, and are added in quadrature.
-See text for future projections
+Table 1: Systematic uncertainty budget from the two-parameter analysis for $b$.All uncertainties are believed to be uncorrelated, and are added in quadrature.
+For details on future projections, see text.
 IV Results
-Our result in the Lee-Yang scalar-tensor Lagrangian of quark-lepton non-SM Lorentz current is then
-$b_{F}$ = -0.0002 $\pm$ 0.024 (stat) $\pm$ 0.039 (syst),
-We explore complementarity of this our result to other experiments next.
+Our result for the Fierz interference term is then $b$ = 0.033 $\pm$ 0.084 (stat) $\pm$ 0.039 (syst). We plot this 1 $\sigma$ constraint, using Eq. 3 only, on Fig. 5 (the thick black lines).
+We also plot (the large black ellipse) our 1 $\sigma$ allowed area considering, as mentioned above, our present work’s measured $A_{\beta}$ and $b$, their correlation through Eqs.
+3–4, our collaboration’s measured 37K $Ft$ value [19], and the average $Ft$ value of $0^{+}\rightarrow 0^{+}$ decays [32].
+This produces a finite constraint on Lorentz tensor couplings and a considerably tighter constraint on Lorentz scalar couplings in Fig.
+5.
+We explore complementarity of our results to other experiments next.
 IV.1 Sensitivity to new physics
-We show a 90% CL (i.e.
-1.65 $\sigma$) exclusion plot of Lorentz scalar vs.
+We show a 1 $\sigma$ exclusion plot of Lorentz scalar vs.
 tensor lepton-quark couplings in Fig.
@@
 Our goal is to show complementarity in situations with less universal physics models than the Lee-Yang Lagrangian.
-Figure 5: Exclusion plot of Lorentz scalar and tensor lepton-quark couplings, 90% CL.
+Figure 5: Exclusion plot at 1 $\sigma$ of Lorentz scalar and tensor lepton-quark couplings to left-handed $\nu$’s and wrong-handed $\beta$’s.
 Consistency between neutron $\beta$ decay, nuclear $\beta$ decay, and particle physics requires a non-universal interaction with scale less than TeV.
-See details and references in text.
-Our 37K results could be said to complement the neutron results on $C_{S}$ because our smaller Gamow-Teller strength weighs $C_{S}$ more than $C_{T}$, yet our own group’s direct constraints on $C_{S}$ from 38mK decay [26] are considerably tighter, shown here averaged with the 32Ar experiment with similar sensitivity [24, 25].
+Our present 37K constraints from $b$ (solid black lines), along with our tighter constraints combining $b$, $A_{\beta}$ and fTfT info (black ellipse), are also detailed in text.
+Our 37K results could be said to complement the neutron results on $C_{S}$, because our smaller Gamow-Teller strength weighs $C_{S}$ more than $C_{T}$.
+Yet our own group’s direct constraints on $C_{S}$ from 38mK decay [27] are tighter, shown here averaged with the 32Ar experiment with similar sensitivity [25, 26].
 We show three constraints from particle physics with computable sensitivity to the Lee-Yang Lagrangian and no sensitivity to 2nd-class currents.
@@
 The other tight indirect constraints on $C_{S}$, shown in the green vertical band, are from the consistency of $0^{+}\rightarrow 0^{+}$ $Ft[E_{\beta}]$ values in a 2020 global survey [32].
-Yet allowing $b_{F}$ to float instead of fixing it to zero would inflate the uncertainty in $V_{ud}$ by factors of two, motivating independent determinations of $b_{F}$ with similar sensitivity.
+Yet allowing $b$ to float instead of fixing it to zero is known to inflate the uncertainty in $V_{ud}$ by about two [14], motivating independent determinations of $b$ with similar sensitivity.
 Recent theory developments may relax or improve the sensitivity: re-evaluation of the lepton phase space $f$ summarized in Ref.
@@
 Two experiments are shown for neutron $\beta$ decay.
-Measurement of $\beta$ decay asymmetry and spectrum shape has competitive sensitivity to $b_{F}$ consistent with zero (solid black lines) [32].
+Measurement of $\beta$ decay asymmetry and spectrum shape has competitive sensitivity to $b$ consistent with zero (solid red lines) [33].
 An analysis including the recoil energy spectrum of protons and Ref.
-[32] data together, but floating $g_{A}/g_{V}$ to a different result from either, asserts a 2.8 $\sigma$ nonzero $b_{F}$ result (the dashed black line) [33].
+[33] data together, but floating $g_{A}/g_{V}$ to a different result from either, asserts a 2.8 $\sigma$ nonzero $b$ result (the dashed red line) [34].
 By inspection, the positive result of Ref.
@@
 Gamow-Teller decay (the cyan band denoted $P_{F}/P_{GT}$[39].
-Neverthless, we note that interpreting the $C_{S}\approx$ -0.05 region as an induced scalar for neutron decay $e\approx$ 31 alleviates tension with all other nuclear $\beta$ decay because of its $1/A$ dependence.
+Neverthless, we note that interpreting the $C_{S}\approx$ -0.05 region as an hypothetical induced scalar for neutron decay with constant form factor $e\approx$ 31 alleviates tension with all other nuclear $\beta$ decay because of the resulting $1/A$ dependence of its contribution.
 We see from Fig.
-5 that to be complementary to other physics, we would have to improve our 37K $b_{F}$ result to compete with the best isospin mirror nucleus results in 19Ne [35, 34, 36]; measurements of $a_{\beta\nu}$ in 21Na [37] would need similar improvement.
-We hope demonstrated and in-progress improvements in the Experiment suggestions will let us achieve the factor of five needed.
+5 that to be complementary to other physics, we would have to improve our 37K $b$ result to compete with the best isospin mirror nucleus results in 19Ne [36, 35, 37] and to other measurements.
+As we are presently dominated by statistics, and project to improve systematics, we hope our demonstrated and in-progress improvements in the experiment will let us achieve the order of magnitude needed.
 V Conclusions
-We have used our $\beta$ asymmetry data from the decay of 37K to constrain new physics that depends on the $\beta$ energy, measuring the Fierz interference term from quark-lepton Lorentz scalar and vector couplings $b_{F}$ = -0.0002 $\pm$ 0.024 (stat) $\pm$ 0.039 (syst), We provide constraints on a combination of Lorentz scalar and tensor lepton-quark couplings.
-Our present results using our laser-cooled atoms are dominated by a combination of statistics and solvable $\beta$ scattering, and we hope to improve sensitivity by a factor of about five to enable complementarity with experiments indicating possible existence of non-standard model physics.
+We have used our $\beta$ asymmetry data from the decay of 37K to constrain new physics that depends on the $\beta$ energy, measuring the Fierz interference term $b$ = 0.033 $\pm$ 0.084 (stat) $\pm$ 0.039 (syst).
+We also provide constraints on an exclusion plot on a combination of Lorentz scalar and tensor lepton-quark couplings from our determination of $b$ and $A_{\beta}$ adding the measured $fT$ of 37K and its uncertainty [19] (using the center value of $fT$ of $0^{+}\rightarrow 0^{+}$ decays for the Fermi strength).
+Our present results using our laser-cooled atoms are dominated by a combination of statistics and solvable $\beta$ scattering, and we hope to improve our statistics-limited result by an order of magnitude to enable complementarity with experiments indicating possible existence of non-standard model physics.
 Acknowledgements.
+We acknowledge physics input from Martín González-Alonso.
 We acknowledge TRIUMF/ISAC staff, in particular for TiC target preparation.